### v2

@@ Abstract
In this work we study the problem of Stochastic Budgeted Multi-round Submodular Maximization (SBMSm), in which we would like to adaptively maximize the sum over multiple rounds of the value of a monotone and submodular objective function defined on a subset of items, subject to the fact that the values of this function depend on the realization of stochastic events and the number of items that we can select over all rounds is limited by a given budget. This problem extends, and generalizes to multiple round settings, well-studied problems such as (adaptive) influence maximization and stochastic probing.
We first show that, if the number of items and stochastic events is somehow bounded, there is a polynomial time dynamic programming algorithm for SBMSm. Then, we provide a simple greedy approximation algorithm for SBMSm, that first non-adaptively allocates the budget to be spent at each round, and then greedily and adaptively maximizes the objective function by using the budget assigned at each round. Such algorithm guarantees a $(1-1/e-\epsilon)$-approximation to the optimal adaptive value.
Finally, by introducing a metric called budget-adaptivity gap, we measure how much an optimal policy for SBMSm, that is adaptive in both the budget allocation and item selection, is better than an optimal partially adaptive policy that, as in our greedy algorithm, determined the budget allocation in advance. We show a tight bound of $e/(e-1)$ on the budget-adaptivity gap, and this result implies that our greedy algorithm guarantees the best approximation among all partially adaptive policies.
@@ 1 Introduction
Consider the well-known problem of influence maximization: here we are given a (social) network, with edge probabilities encoding the strength of the relationships among network’s components; these probabilities describe how easily influence (e.g., an information or an innovation) spreads over the network starting from a limited set of nodes, usually termed seeds; the goal is then to find the set of seeds that maximizes the expected number of “influenced” nodes. The non-adaptive version of the problem, in which all the seeds are chosen before that influence starts, has been widely studied, with numerous successful algorithms [32, 12, 8] and applications ranging from viral marketing [32], to election manipulation [39, 16, 10], campaigns to contrast misinformation [2] to infective disease prevention [40]. Recently, focus moved to the adaptive variant of this problem (in which seeds can be added even when the influence is spreading from previously placed seeds) [26, 34, 17]. This variant happens to have strong relationship with some Stochastic Probing problems [4, 3], in which there are $n$ items whose states are stochastically defined according to (not necessarily equal) independent distributions and the goal is to choose a feasible subset of items maximizing the expected value of a non-negative, monotone and submodular set function. Several stochastic probing settings can be also encountered in expert hiring [30], online dating and kidney exchange [27, 9], Bayesian mechanism design and sequentially posted pricing [11, 1].
@@
Research both in influence maximization and stochastic probing focused almost exclusively on the single-round setting. Recently, there is an increasing interest in those cases in which the optimization is run in a multi-round setting. This is highlighted both by the increasing amount of literature on the topic (see Related Works), and by the numerous applications that are evidently deployed over multiple rounds, as it is the case, e.g., for many real-word marketing, electoral and misinformation/disinformation contrasting campaigns. Moreover, it is often the case that in these contexts we are interested both in maximizing the total effectiveness of the campaign, and in limiting the total budget spent. Note that, in order to achieve this goal it may be insufficient to limit the budget spent at each round, since this may lead in general to very inefficient outcomes (see Remark 1 below for more details). However, this makes the problem much more complex: we do not only need to care about which actions to take at each round, but also we need to care about how large must be the portion of the budget spent at that round.
@@
Unfortunately, $\text{SRm}(t,b)$ turns out to usually be a hard problem to solve. Indeed, it can be easily reduced from the (deterministic) monotone submodular maximization problem, that is known to be APX-hard [23], and there is no algorithm that approximates this problem in general instances within a factor better than $1-1/e$ (unless P=NP). Nevertheless, we show that there are non-trivial instances for which the problem can be effectively solved in polynomial time, essentially corresponding to the cases in which there are a few items and the possible states of the underlying stochastic environment are polynomial in the size of the input instance. The obtained findings imply that both $\text{SRm}(t,b)$ and SBMSm are fixed-parameter-tractable (FTP) problems [22] w.r.t. to the number of items, if the outcomes of the stochastic environment are polynomially bounded.
@@
For this reason, we need to resort to a different framework. Interestingly, we will show that a greedy polynomial time algorithm, that is arguably simpler than the previous one, provides a $(1-1/e-\epsilon)$-approximation of the optimal adaptive solution for every instance, where $\epsilon>0$ is an arbitrarily small constant due to the use of some randomized sampling procedures. Hence, the obtained approximation matches that guaranteed in the much simpler single-round case by [26] and cannot be substantially improved in light of the hardness result of [23]; we also point that the machinery of [26] cannot be directly applied to our multi-round problem to obtain the same approximation guaranteed for the restricted single-round setting (see Remark 2 below), and this further made the introduction of a new and more general algorithmic approach necessary.
Interestingly, our greedy approximation algorithm relies on a result that can be of independent interest. Indeed, we prove that in a SBMSs problem, a non-adaptive allocation of budget to rounds can be sufficient to guarantee a constant approximation of the optimal adaptive solution. Even more surprisingly, we can achieve such result by allocating the budget to rounds greedily, i.e., by assigning each unit of budget to the round in which it is expected to produce the largest expected value of the objective function, given the amount of budget still available. Our ($1-1/e-\epsilon)$-approximate algorithm then follows by composing this greedy allocation of budget, with the algorithm that chooses the items to observe up to the allocated budget in a greedy adaptive way, i.e., it chooses an item that is expected to achieve the approximate maximum value of the objective function with high probability (with both approximation and probability depending on the choice of $\epsilon>0$), observes the realization of the state for this item, and then it chooses the next item among the ones that are expected to maximize the objective function given the state realization for the first item, and so on. In other words, our result shows that the powerful optimal adaptive solution for an SBMSs problem, that is adaptive both in the allocation of budget to rounds and in the choice of selected items within rounds, can be well-approximated by an algorithm that is only partially adaptive, i.e., it is adaptive in the choice of selected items, but not in the allocation of budget to rounds.
@@
We show that the budget-adaptivity gap is exactly $e/(e-1)\approx 1.582$, that is, the non-adaptive budget allocation, in general, cannot be used to reach the fully adaptive optimum but still constitutes a good approximation of the fully adaptive ones; this result also implies that the above greedy algorithm essentially guarantees the best approximation among all partially adaptive policies. We point out that the effectiveness of a partially adaptive approach for multi-round adaptive optimization has been left as prominent research direction by [6], and to the best of our knowledge, our results are the first that consider such an approach and guarantee, at the same time, a good approximation to the optimal solution in polynomial time.
@@ 2 Model and Definitions
Given two integers $k_{1},k_{2}$, let $\llbracket k_{1},k_{2}\rrbracket:=\{k_{1},\ldots,k_{2}\}$ be the discrete interval from $k_{1}$ to $k_{2}$ (that is empty if $k_{1}>k_{2}$); furthermore, let $[k]$ denote the discrete interval $\llbracket 1,k\rrbracket$. Given an event $A$, let $\bm{1}_{A}$ denote the boolean variable that is equal to $1$ if $A$ is true/happens, and equal to 00 otherwise.
@@ 2.3 Properties of the Objective Functions
Henceforth, we assume that objective functions satisfy all the properties described above.
@@ 2.5.1 Influence Maximization
Mapping between IM and SBMSm.
  • $T$ and $B$ are exactly the same as in $M$ (hence, we use exactly the same name).
  • The set $V$ of items is the set of nodes in the social network.
Properties of $f_{t}$.
@@ 2.5.2 Stochastic Probing
Mapping between MR-BSS and SBMSm.
Properties of $f_{t}$.
@@ 3.2 How to Solve Single-round Optimization Efficiently?
In the above dynamic-programming algorithm, we assumed the existence of a sub-routine that solves $\text{SRm}(t,b)$ for any round $t$ and budget $b$, i.e., that allows to compute recurrence (2). Unfortunately, the $\text{SRm}(t,b)$ problem can be easily reduced from the submodular monotone optimization problem, that is NP-hard [23]. To deal with the intractability of the single-round optimization problem, we will describe an algorithm that is in general inefficient, but that can be run in polynomial time in a few cases.
The Single-round Optimization Problem as a Game.
From the Game to the Optimal Solution.
When the Single-round Problem can be Efficiently Solved?
Proposition 1.
${\cal T}_{t}$ has $O(2^{n}n!|H|)$ nodes.
Proof.
$|{\cal T}_{t}|\in O(z).$ (3)
Now, we show that
$z\in O(2^{n}n!|H|).$ (4)
Corollary 1.
$\text{SRm}(t,b)$ can be solved in $O(2^{n}n!|H|)$ time.
Corollary 2.
SBMSm can be solved in $O(2^{n}n!nT^{2}|H|)$ time.
Finally, by exploiting again the complexity bounds provided in Corollaries 1 and 2, we have the following result on fixed-parameter-tractability:
Corollary 3.
Both $\text{SRm}(t,b)$ and SBMSb are fixed-parameter-tractable problems w.r.t. to $n$, if $|H|$ is polynomially bounded in the size of the input instance.
What Happens if the Single-round Problem cannot be Efficiently Solved?
@@
While it would be possible to fix our dynamic programming framework in order to do not cause such an explosion in the approximation guarantee, and to provide an $\left(1-1/e-\epsilon\right)$-approximation to the optimal multi-round policy, we next prove that the same bound can be actually achieved by a simpler and less computationally expensive greedy algorithm.
@@ 4 Partially Adaptive Approximation Algorithm
In this section, we will design and analyze a partially adaptive policy as approximation algorithm for SBMSm. In general, a partially adaptive policy, in a first part, determines how to distribute the budget $B$ over all rounds by computing a budget vector $\vec{b}=(b_{1},\ldots,b_{T})$, so that $b_{t}$ is the budget assigned to round $t$; this budget assignment is done in a non-adaptive way, that is, without knowing the particular realization of the stochastic environment, but only based on the knowledge of the input instance $I$. Subsequently, for each round $t\in[T]$, the policy adaptively selects $b_{t}$ items to be observed in round $t$; differently from the the first part, now the realization of the stochastic environment is partially revealed after each observation. It is worth to notice that an optimal solution for SBMSm, in general, dynamically allocates the budget after each item selection. Instead a partially adaptive policy allocates all the budget to the rounds at the beginning.
The specific partially adaptive policy we will consider is based on the greedy approach: (i) it first determines the budget allocation by repeatedly assigning each unit of budget to the rounds in a non-adaptive greedy way, and then (ii) greedily selects the items in an adaptive way (according to the budget previously allocated to each round). Despite partially adaptive policies are in general weaker than fully adaptive ones (because of the relaxation on budget adaptivity), we will show that our policy still guarantees a good approximation of the optimal adaptive value. In particular, we will show that its approximation factor is $1-1/e-\epsilon\approx 0.632-\epsilon$, where $e\approx 2.718$ denotes the Nepero number and $\epsilon>0$ is an arbitrarily small error due to some sampling procedures. Furthermore, as a consequence of the hardness result of [23] (holding for non-adaptive single-round submodular maximization), we have that SBMSm cannot approximated within a factor better than $1-1/e$ (unless P=NP). Thus, the approximation guaranteed by our partially adaptive policy cannot be substantially improved.
@@ 4.1 Estimating Objective Functions
In order to solve the desired maximization problems through our greedy partially adaptive policy, we first need to estimate the expected value of the objective functions achieved when selecting an item. However, this is not an easy task due to the stochastic nature of the problem. We first introduce a standard Monte-Carlo algorithm, called ${\sf Oracle1}$, that can be used to give a good estimation, with high probability, of the expected value of the objective function when selecting an item: such algorithm can be used to find an item that approximately maximizes the expected influence spread of an objective function, under a given partial state.
@@ 4.1.1 Oracle1
Given $\delta,\xi>0$, ${\sf Oracle1}$ satisfies the following property.
@@
Proof.
$\leq n\cdot 2\exp\left(-\frac{2(\delta/2)^{2}q}{\Lambda^{2}}\right)$
$=\xi.$
$\mathbb{P}\left[|\Delta(v|\langle S_{t}\rangle)-\tilde{\Delta}(v|\langle S_{t}\rangle)|\leq\delta/2\ \forall v\in V\right]$ $=\mathbb{P}\left[\Bigg{|}\mathbb{E}\left[\bm{X}_{t,v,\langle S\rangle}\right]-\sum_{j=1}^{q}\frac{X^{j}_{t,v,\langle S\rangle}}{q}\Bigg{|}\leq\delta/2\ \forall v\in V\right]$ (5)
$\geq 1-\xi.$
$\mathbb{P}\left[\max_{v\in V}\Delta(v|\langle S_{t}\rangle)-{\Delta}(\tilde{v}|\langle S_{t}\rangle)\leq\delta\right]$ $=\mathbb{P}\left[\Delta(v^{*}|\langle S_{t}\rangle)-{\Delta}(\tilde{v}|\langle S_{t}\rangle)\leq\delta\right]$
$\geq\mathbb{P}\left[(\Delta(v^{*}|\langle S_{t}\rangle)-\tilde{\Delta}(v^{*}|\langle S_{t}\rangle))+(\tilde{\Delta}(\tilde{v}|\langle S_{t}\rangle)-{\Delta}(\tilde{v}|\langle S_{t}\rangle))\leq\delta\right]$ (6)
$\geq\mathbb{P}\left[|\Delta(v|\langle S_{t}\rangle)-\tilde{\Delta}(v|\langle S_{t}\rangle)|\leq\delta/2\ \forall v\in\{\tilde{v},v^{*}\}\right]$
$\geq 1-\xi,$ (8)
where (6) holds as $-\tilde{\Delta}(v^{*}|\langle S_{t}\rangle)+\tilde{\Delta}(\tilde{v}|\langle S_{t}\rangle)\geq 0$ (by the optimality of $\tilde{v}$ for $\tilde{\Delta}(\tilde{v}|\langle S_{t}\rangle)$), (7) holds since $\Delta(v|\langle S_{t}\rangle)=\mathbb{E}\left[\bm{X}_{t,v,\langle S_{t}\rangle}\right]$ and $\tilde{\Delta}(v|\langle S_{t}\rangle)=\sum_{j=1}^{q}\frac{X^{j}_{t,v,\langle S\rangle}}{q}$, and (8) follows from (5). Finally, Property (i) follows from (5).
Property (ii) immediately follows from: (a) the polynomial boundedness property, that polynomially relates the quantity $\Lambda$ (appearing in $q(\delta,\xi)$) to the size of the input instance; (b) the existence of an oracle that picks samples drawn from the underlying probability distributions (whose existence has been implicitly assumed in the model definition of Section 2). ∎

### v3

@@ Abstract
In this work, we study the Stochastic Budgeted Multi-round Submodular Maximization (SBMSm) problem, where we aim to adaptively maximize the sum, over multiple rounds, of a monotone and submodular objective function defined on subsets of items. The objective function also depends on the realization of stochastic events, and the total number of items we can select over all rounds is bounded by a limited budget. This problem extends, and generalizes to multiple round settings, well-studied problems such as (adaptive) influence maximization and stochastic probing.
We show that, if the number of items and stochastic events is somehow bounded, there is a polynomial time dynamic programming algorithm for SBMSm. Then, we provide a simple greedy $1/2(1-1/e-\epsilon)\approx 0.316$-approximation algorithm for SBMSm, that first non-adaptively allocates the budget to be spent at each round, and then greedily and adaptively maximizes the objective function by using the budget assigned at each round. Finally, we introduce the budget-adaptivity gap, by which we measure how much an adaptive policy for SBMSm is better than an optimal partially adaptive one that, as in our greedy algorithm, determines the budget allocation in advance. We show that the budget-adaptivity gap lies between $e/(e-1)\approx 1.582$ and $2$.
@@ 1 Introduction
Consider the well-known problem of influence maximization: here we are given a (social) network, with edge probabilities encoding the strength of the relationships among network’s components; these probabilities describe how easily influence (e.g., an information or an innovation) spreads over the network starting from a limited set of nodes, usually termed seeds; the goal is then to find the set of seeds that maximizes the expected number of “influenced” nodes. The non-adaptive version of the problem, in which all the seeds are chosen before that influence starts, has been widely studied, with numerous successful algorithms [33, 13, 9] and applications ranging from viral marketing [33], to election manipulation [42, 17, 11], campaigns to contrast misinformation [3] to infective disease prevention [43]. Recently, focus moved to the adaptive variant of this problem (in which seeds can be added even when the influence is spreading from previously placed seeds) [27, 37, 19]. This variant happens to have strong relationship with some Stochastic Probing problems [5, 4], in which there are $n$ items whose states are stochastically defined according to (not necessarily equal) independent distributions and the goal is to choose a feasible subset of items maximizing the expected value of a non-negative, monotone and submodular set function. Stochastic probing is relevant in various applications, such as the optimal placement of sensors for environmental monitoring [34] or energy resources within electrical distribution networks [1, 36]. In these scenarios, items represent potential locations within a network, where each item’s state models the (uncertain) efficiency of a sensor or energy resource placed in the corresponding location. Feasible solutions correspond to sets of locations where sensors or resources can be deployed and/or activated, subject to cost or pollution constraints. The submodular set function models either the coverage provided by the sensors or the energy efficiency achieved. Other stochastic probing settings can be also encountered in expert hiring [31], online dating and kidney exchange [28, 10], Bayesian mechanism design and sequentially posted pricing [12, 2].
@@
Research both in influence maximization and stochastic probing focused almost exclusively on the single-round setting. Recently, there is an increasing interest in those cases in which the optimization is run in a multi-round setting. This is highlighted both by the increasing amount of literature on the topic (see Related Works), and by the numerous applications that are evidently deployed over multiple rounds, as it is the case, e.g., for many real-word marketing, electoral and misinformation/disinformation contrasting campaigns. Indeed, it is often the case that in these contexts we are interested both in maximizing the total effectiveness of the campaign, and in limiting the total budget spent. Note that, in order to achieve this goal it may be insufficient to limit the budget spent at each round, since this may lead in general to very inefficient outcomes (see Remark 1 below for more details). However, this makes the problem much more complex: we do not only need to care about which actions to take at each round, but also we need to care about how large must be the portion of the budget spent at that round. Similarly, in contexts involving the optimal use of sensors or energy resources over time, it’s possible that the cost or pollution constraints are not imposed daily, but rather on an average basis over the entire time-horizon. This means that while certain days may exceed these limits, the overall usage throughout the planning period must meet the constraints on average.
@@
Unfortunately, $\text{SRm}(t,b)$ turns out to usually be a hard problem to solve. Indeed, it can be easily reduced from the (deterministic) monotone submodular maximization problem, that is known to be APX-hard [24], and there is no algorithm that approximates this problem in general instances within a factor better than $1-1/e$ (unless P=NP). Nevertheless, we show that there are non-trivial instances for which the problem can be effectively solved in polynomial time, essentially corresponding to the cases in which there are a few items and the possible states of the underlying stochastic environment are polynomial in the size of the input instance. The obtained findings imply that both $\text{SRm}(t,b)$ and SBMSm are fixed-parameter tractable (FTP) problems [23] w.r.t. to the number of items, if the outcomes of the stochastic environment are polynomially bounded.
@@
For this reason, we need to resort to a different framework. Interestingly, we will show that a greedy polynomial time algorithm, that is arguably simpler than the previous one, provides a $1/2(1-1/e-\epsilon)\approx 0.316$-approximation of the optimal adaptive solution for every instance, where $\epsilon>0$ is an arbitrarily small constant due to the use of some randomized sampling procedures. Even more surprisingly, we can achieve such result in two phases: (i) we first preallocate the budget to rounds greedily, i.e., by assigning each unit of budget to the round in which it is expected to produce the largest expected value of the objective function, given the amount of budget still available; (ii) then, we apply the adaptive-greedy algorithm [27] that chooses, at each round, the items to observe up to the allocated budget in a greedy adaptive way, i.e., it sequentially chooses an item that is expected to achieve the approximate maximum value of the objective function with high probability (with both approximation and probability depending on the choice of $\epsilon>0$), observes the realization of the state for this item, and then it chooses the next item among the ones that are expected to maximize the objective function given the state realization for the first item, and so on. To show the $1/2(1-1/e-\epsilon)$-approximation factor, we first prove that in a SBMSs problem there exists a non-adaptive allocation of budget to rounds that guarantees a $1/2$-approximation of the optimal adaptive solution. Then, we show that the adaptive greedy algorithm (considered in the phase (ii) described above) applied to such budget allocation further lowers the approximation guarantee by a factor of $(1-1/e-\epsilon)$. Finally, we show that the best non-adaptive allocation of budget for the adaptive greedy algorithm is that determined by preallocating the budget to rounds greedily, as in the phase (i) described above.
In other words, our result shows that the powerful optimal adaptive solution for an SBMSs problem, that is adaptive both in the allocation of budget to rounds and in the choice of selected items within rounds, can be well-approximated by an algorithm that is only partially adaptive, i.e., it is adaptive in the choice of selected items, but not in the allocation of budget to rounds.
@@
We show that the budget-adaptivity gap is between $e/(e-1)\approx 1.582$ and $2$, that is, the non-adaptive budget allocation, in general, cannot be used to reach the fully adaptive optimum but still constitutes a good approximation of the fully adaptive ones; this result also implies that the above greedy algorithm essentially guarantees the best approximation among all partially adaptive policies. We point out that the effectiveness of a partially adaptive approach for multi-round adaptive optimization has been left as prominent research direction by [7], and to the best of our knowledge, our results are the first that consider such an approach and guarantee, at the same time, a good approximation to the optimal solution in polynomial time.
@@ 2 Model and Definitions
Given two integers $k_{1},k_{2}$, let $\llbracket k_{1},k_{2}\rrbracket:=\{k_{1},\ldots,k_{2}\}$ be the discrete interval from $k_{1}$ to $k_{2}$ (that is empty if $k_{1}>k_{2}$); furthermore, let $[k]$ denote the discrete interval $\llbracket 1,k\rrbracket$.
@@ 2.3 Properties of the Objective Functions
In this work, we strengthen the adaptive submodularity by introducing the strong adaptive submodularity, that is defined as follows.
Definition 3.
An objective function $f_{t}:2^{V}\rightarrow{\mathbb{R}}_{\geq 0}$ is strongly adaptive submodular if the following properties hold: (i) It is adaptive submodular. (ii) Given an integer $k\geq 1$, let $\pi$ be an optimal adaptive policy that selects $k$ items. Then, for any partial state $\langle S_{t}\rangle$ and any adaptive selection $\tilde{\bm{S}}_{t}\supseteq S_{t}$ of at most $k+|S_{t}|$ items conditioned by the observation of $\langle S_{t}\rangle$, we have
$\sigma(\pi)\geq\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\tilde{\bm{S}}_{t},\bm{\eta}_{t})-f_{t}(S_{t},\bm{\eta}_{t})|\langle S_{t}\rangle\prec\bm{\eta}_{t}].$ (1)
Intuitively, adaptive submodularity (i.e., property (i) of strong adaptive submodularity) requires that the marginal increment of the expected value of $f_{t}$ due to the selection of the same item $v$ does not increase as the observations become richer. In contrast, the second property of strong adaptive submodularity requires that the expected increment of $f_{t}$, achieved by the adaptive selection of $k$ items after observing a given partial state, is not better than the optimal adaptive value for selecting $k$ items, without any prior item selection and observation.
From this point forward, we assume that the objective functions satisfy all the properties described above.
@@ 2.5.1 Influence Maximization
In Appendix A we show that MR-IM-IC can be modeled as an SBMSm problem satisfying all required properties.
@@ 2.5.2 Stochastic Probing
In Appendix A we show that MR-BSS can be modeled as an SBMSm problem satisfying all required properties.
@@ 3 Optimal Algorithm
We stress that the properties given in Section 2.3 are not required to hold in the results provided in this section (differently from the subsequent sections).
@@ 3.2 How to Solve Single-round Optimization Efficiently?
In the above dynamic-programming algorithm, we assumed the existence of a sub-routine that solves $\text{SRm}(t,b)$ for any round $t$ and budget $b$, i.e., that allows to compute recurrence (3). Unfortunately, the $\text{SRm}(t,b)$ problem can be easily reduced from the submodular monotone optimization problem, that is NP-hard [24]. To address the intractability of the single-round optimization problem, in Appendix B we describe a tree-based algorithm that, while generally inefficient, can be run in polynomial time in specific cases. In particular, we show that this algorithm is FPT with respect to $n$, provided the number of states $|H|$ is polynomially bounded in the input instance. Consequently, by incorporating this algorithm as a subroutine in the dynamic programming approach described above, we obtain an FPT algorithm for the more general SBMSm problem, as outlined in the following theorem.
Theorem 1 (Corollary 3 of Appendix B).
Both $\text{SRm}(t,b)$ and SBMSm are FPT problems w.r.t. to $n$, if $|H|$ is polynomially bounded in the size of the input instance.
@@ 4 Partially Adaptive Approximation Algorithm
In this section, we will design and analyze a partially adaptive policy as $1/2(1-1/e-\epsilon)\approx 0.316$-approximation algorithm for SBMSm, where $e\approx 2.718$ denotes the Nepero number and $\epsilon>0$ is an arbitrarily small error due to some sampling procedures. This policy is based on the greedy approach, and its approximation guarantee and time-complexity hold under the properties given in Section 2.3.
In general, a partially adaptive policy, in a first part, determines how to distribute the budget $B$ over all rounds by computing a budget vector $\vec{b}=(b_{1},\ldots,b_{T})$, so that $b_{t}$ is the budget assigned to round $t$; this budget assignment is done in a non-adaptive way, that is, without knowing the particular realization of the stochastic environment, but only based on the knowledge of the input instance $I$. Subsequently, for each round $t\in[T]$, the policy adaptively selects $b_{t}$ items to be observed in round $t$; differently from the the first part, now the realization of the stochastic environment is partially revealed after each observation. It is worth to notice that an optimal solution for SBMSm, in general, dynamically allocates the budget after each item selection. Instead a partially adaptive policy allocates all the budget to the rounds at the beginning.
The specific partially adaptive policy we will consider is based on the greedy approach: (i) it first determines the budget allocation by repeatedly assigning each unit of budget to the rounds in a non-adaptive greedy way, and then (ii) greedily selects the items in an adaptive way (according to the budget previously allocated to each round).
Despite partially adaptive policies are in general weaker than fully adaptive ones (because of the relaxation on budget adaptivity), we will show that our policy still guarantees a good approximation of the optimal adaptive value. Indeed, as a consequence of the hardness result of [24] (holding for non-adaptive single-round submodular maximization), we have that SBMSm cannot be approximated within a factor better than $0.632\approx 1-1/e$ (unless P=NP). Thus, the approximation guaranteed by our partially adaptive policy is close to the known inapproximability bound.
@@ 4.1 Estimating Objective Functions
In order to solve the desired maximization problems through our greedy partially adaptive policy, we first need to estimation the expected value of the objective functions achieved when selecting an item. However, this is not an easy task due to the stochastic nature of the problem. We first introduce a standard Monte-Carlo algorithm, called ${\sf Oracle1}$, that can be used to give a good estimation, with high probability, of the expected value of the objective function when selecting an item: such algorithm can be used to find an item that approximately maximizes the expected influence spread of an objective function, under a given partial state.
@@ 4.1.1 Oracle1
Given $\delta,\xi>0$, ${\sf Oracle1}$ satisfies the following property, whose proof is deferred to Appendix C.